OpenStudy (laddiusmaximus):

calculate the derivative. tln(t^(2) +1)

OpenStudy (kinggeorge):

This requires application of the product rule and the chain rule. First apply the product rule, and then apply the chain rule. You should get \[{d \over dt}\;\;t \ln(t^2+1)=1*\ln(t^2+1) + t*{1\over t^2+1} *2t\]I trust that you can simplify any further.